OpenStudy (anonymous):

Solve x2 - 8x + 3 = 0 by completing the square. Which equation is used in the process? (x - 4)^2 = 0 (x - 4)^2 = 13 (x - 4)^2 = 19

OpenStudy (anonymous):

@xoya

OpenStudy (anonymous):

The second option

OpenStudy (anonymous):

x2-8x+3=0 x2-2*4*x+16-13=0 (x-4)^2=13 so, it's (b)

OpenStudy (anonymous):

LOL yeahh i was right :D

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

x^2-8x+3=0 Here we are using this formula (a-b)^2=a^2-2ab+b^2 x^2-2(x)(4)+(4)^2-13=0 (x-4)^2=0+13 (x-4)^2=13

OpenStudy (anonymous):

Your welcome :D